# Bottom-up Analysis of Ro-Vibrational Helical Dichroism

### v1

@@ Abstract
Helical dichroism (HD) is a proposed method for the resolution of molecular chirality, employing the orbital angular momentum (OAM) of light. Going beyond the conventional assumptions about HD, this work proposes a rigid theoretical framework for the analysis of the HD, based on molecular symmetries and rotational eigenstates. We derive the rotational selection rules, which clearly establish that HD only emerges from the spin-orbit coupling of light, even for beams without far-field OAM. Our findings refine the conditions for observing HD, shedding light on the outcome of prior experiments and guiding future designs for chiral sensing using structured light.
@@
For decades, the difference in absorption of left- and right-circularly polarised light, known as the circular dichroism (CD) [6, 7], has been used in spectroscopy of chiral media [8]. However, circular dichroism relies on a limited resource, namely the spin angular momentum of light. Meanwhile, pioneering optics work by Allen et al. [9] has demonstrated that, in addition to spin, light also carries orbital angular momentum (OAM). In contrast to spin, the number of OAM quanta a beam can carry is, in principle, unbounded. The helical wavefronts of OAM-carrying light fascinated physicists, who envisaged that their chiral structure could match the chirality of matter, offering unprecedented possibilities for enantioselective applications.
@@
This fascination culminated in concrete examples of OAM-induced processes being proposed as probes of enantiomeric purity, particularly in the form of helical dichroism (HD) [10, 11, 12], which is argued to represent a significant improvement over CD [11, 12]. The seminal theoretical works of Forbes and Andrews [13, 14, 15] laid the groundwork for HD, carefully analyzing the symmetry properties of the terms in the multipolar expansion of light-matter interaction and pinpointing the electric-dipole–magnetic-dipole ($E_{1}$-$M_{1}$), the electric-dipole–electric-quadrupole ($E_{1}$-$E_{2}$) and electric-quadrupole–electric-quadrupole ($E_{2}$-$E_{2}$) terms as chirality discriminators. They derived the expression for the $E_{1}$-$E_{2}$ transition amplitude in terms of molecular multipole moments; however, they made no attempt to evaluate the latter in terms of molecular rotational states. On the experimental side, initial attempts showed no signs of HD [16, 17]; however, recent experiments have claimed to observe it [18, 19, 20, 21, 22]. Nonetheless, the experimental apparatus of these experiments is very complex, and, although it was argued that tight focusing is a requirement [14], there have so far been no theoretical attempts to pinpoint the origin of the observed dichroism.
@@
This work fills this gap in the literature by unravelling the conditions under which HD can arise in ro-vibrational spectroscopy. We first provide an overview of the chirality detection methods based on molecular multipole moments and derive the symmetry requirements for observing the $E_{1}$-$E_{2}$-based dichroism. Using our recently developed molecule-light interaction Hamiltonian [23, 24] we derive the selection rules for ro-vibrational transitions induced by the $E_{1}$-$E_{2}$ term. By analyzing these rules, we demonstrate that paraxial Laguerre-Gaussian (LG) beams cannot produce genuine HD, i.e. HD arising from OAM transfer. In particular, strong focusing is found to be essential, since spin–orbit coupling (SOC) of light [25] enables OAM transfer to chiral molecules, leading to HD contributions that are absent within the paraxial approximation. Furthermore, we demonstrate that even a Gaussian beam, which carries no OAM in the far-field, can produce a HD signal in the tightly-focused regime. Collectively, our findings refine the theoretical foundations of HD in ro-vibrational spectroscopy and guide the designs of future experiments and enantiomer-selective techniques using structured light.
@@
Several existing enantiomer-sensitive methods rely on the analysis of the multipole moments of the molecule. The (vibrational) circular dichroism – the standard method to detect chirality – concerns the $E_{1}$-$M_{1}$ term and has already been exhaustively investigated [26, 27, 28, 29]. It draws on the fact that the electric dipole moment of a chiral molecule transforms as a vector, and thus picks up a minus sign, under mirror reflection, while the magnetic dipole, which is a pseudovector, is invariant to it. A coexistence of both these multipoles implies that a molecule lacks a mirror symmetry plane (strictly speaking an improper rotation axis) and is, by definition, chiral. For this condition to be applicable, the molecular symmetry must permit at least one nonzero dipole moment, enabling this type of CD to probe chiral molecules across several symmetry groups. However, the dichroism stemming from the $E_{1}$-$M_{1}$ term only depends on the intensity of the field, which makes other sources, where the dichroism can be enhanced by the structure of the field, attractive alternatives.
@@
One example, the phase sensitive microwave three-wave mixing technique [30, 31], exploits the fact that a non-zero product of all dipole moments $d_{x}d_{y}d_{z}\neq 0$ implies that the molecule is chiral [30, 32, 33]. Symmetry analysis reveals that only completely asymmetric ($C_{1}$) chiral molecules can be characterized this condition.
@@
Our work explores another direction – the conditions for chirality stemming from the electric dipole and quadrupole moments of the molecule. The $E_{1}$-$E_{2}$ term-induced dichroism can be enhanced with appropriately structured fields, which makes it a promising prospect for future enantiomer-sensitive techniques. Recently, the magnetic-dipole-magnetic quadrupole ($M_{1}$-$M_{2}$) coupling has also been suggested as a source of HD, which is equivalent to the $E_{1}$-$E_{2}$ term from a symmetry standpoint [34]. However, this and other higher-order terms, such as $E_{2}$-$E_{2}$, are difficult to detect experimentally due to low transition probabilities in the multipole expansion.
@@
The remainder of this paper focuses on characterizing the dichroism originating from the $E_{1}$-$E_{2}$ term. Within a simple point-charge model we derive the following condition (proof in the Supplementary material  [^1] See Supplementary Material, containing references [23, 24, 40, 25, 37]):
@@
If the product of dipole and quadrupole moments $d_{\mu}Q_{\nu\rho}\neq 0$ for distinct $\mu,\nu,\rho\in\{x,y,z\}$ coordinates of the principal axis system, then the molecule is chiral.
@@
The above condition has so far always been connected to HD [10, 13]. However, we show in this work, that it can also lead to discriminatory effects without any transfer of orbital angular momentum from the light to the internal molecular degrees of freedom. Within this work HD will be used to denote dichroism effects where such a transfer occurs, while dichroism where no OAM is transferred, even with an OAM-carrying beam is referred to as CD.
@@
Table 1 summarizes for which chiral point groups dichroism induced by the $E_{1}$-$E_{2}$ term can be observed. In the case of pure-rotational spectroscopy, only the permanent multipole moments of the molecule are relevant. In ro-vibrational spectroscopy the vibrational transitions can alter the symmetry of the molecular state. Therefore, the transformation properties of the final vibrational states in terms of their associated irreducible representation $V$ need to be considered, which allows for the detection of more chiral groups. Table 1 provides the list of $V$’s exhibiting dichroism for each chiral point group.
@@
Table 1: The $E_{1}$-$E_{2}$ chiral signatures in rotational and ro-vibrational spectroscopy of all possible chiral point groups. In the case of ro-vibrational spectroscopy we provide the irreducible-representations of the vibrational modes associated with non-zero chiral signature.
Point group $C_{1}$ $C_{2}$ $C_{n\geq 3}$ $D_{2}$ $D_{n\geq 3}$ $T$ $O$ $I$
@@
Ro-vibrational $A$ $A$, $B$ $E_{1}$ $B_{1,2,3}$ $E_{1}$ $T$
As displayed in Table 1 the $E_{1}$-$E_{2}$ term can only identify chiral molecules with $C_{1}$ and $C_{2}$ symmetries via transitions that preserve the molecular symmetry (pure rotational and ro-vibrational of the trivial irrep. $A$). Since most known chiral molecules are $C_{1}$-symmetric, i.e., they are asymmetric tops [36], we restrict our scope to these molecules.
The eigenstates of the asymmetric top rotor Hamiltonian, expressed in the rotational basis |$\ket{J,M,K}$ [36], read
|$\ket{J_{N_{p},N_{o}}(M)}=\sum_{K}C_{K}\ket{J,M,K}.$ (1)
Here, $J$ is total angular momentum quantum number, $M$ and $K$ are the projections of ${\bm{J}}$ on the lab quantization axis and the body-fixed principal axis respectively. The asymetric top rotor indices $N_{p}$ and $N_{o}$ are the limits of $K$ in the pure prolate/oblate solutions and are not good quantum numbers. Since quantum numbers $J$ and $M$ are not affected by the change in basis (1), we proceed to derive the selection rules for the electric dipole ($E_{1}$) and quadrupole ($E_{2}$) transitions within the rotational basis, $|J,M,K\rangle$.
The transition rate from an initial $i=\ket{J,M,K}$ to a final $f=\ket{J,M^{\prime},K^{\prime}}$ molecular rotational state can be calculated using the light-matter interaction Hamiltonian, derived in Ref. [23, 24]. The transition amplitude reads
@@
where $l=0$ and $l=1$ for dipole and quadrupole transitions respectively, $m=0,\pm 1,\dots,\pm l$ is the OAM transferred to the molecule, $\mu=0,\pm 1,\dots,\pm(l+1)$ runs over all components of the chosen multipole moment in the spherical basis, and $\sigma=0,\pm 1$ denotes the polarization component of light. The vibrational, $\mathcal{I}^{\text{vib};l,\mu}$, and center-of-mass, $\mathcal{I}^{\text{CM};l,m}$, terms are parameterizing the molecular and light properties. In this work, we are interested in the rotational transition amplitude $\mathcal{I}^{\text{rot};l,m,\mu,\sigma}_{J,M,K,J,M^{\prime},K^{\prime}}$, which governs the angular momentum selection rules, summarized in Table 2.
@@
|$\ket{J,0,K}\not\to\ket{J,0,K^{\prime}}$ for details)
Table 2: Selection rules for electric dipole ($E_{1}$) and quadrupole ($E_{2}$) transitions between rotational states |$\ket{J,M,K}$. For details on prohibited transitions see [35].
It can be shown that $\mathcal{I}^{\rm rot;l,m,\mu,\sigma}_{J,M,K,J^{\prime},M^{\prime},K^{\prime}}\propto C^{J^{\prime},M^{\prime}}_{l+1,m-\sigma;J,M}$ [23, 24], with $C^{J,M}_{j_{1},m_{1},j_{2},m_{2}}$ the Clebsh-Gordan coefficient (derivation in the Supplementary material [35]). To realize an $E_{1}$-$E_{2}$ transition both $E_{1}$ ($l=m=0$) and $E_{2}$ ($l=1$, $m=0,\pm 1$) transitions should be allowed between the chosen states simultaneously. For light with a well-defined polarization, $\sigma$, this is only possible for the $m=0$ component of the beam, which carries no OAM (see Table 2). This brings us to a first key finding of our work: The dichroism stemming from the $E_{1}$-$E_{2}$ term of a paraxial beam should not be associated with HD, as no OAM transfer to the molecule occurs. Moreover, this dichroism is present even in the case of a plane-wave and thus is not associated at all to the structure of light or helicity.
This raises the question whether the $E_{1}$-$E_{2}$ term can produce HD, where OAM transfer occurs. We show in what follows that we can make such a construction in the case of non-paraxial beams.
In Fig. 1, we demonstrate the allowed dipole (red squares) and quadrupole (blue squares) transitions between the rotational eigenstates of asymmetric top molecules. For clarity, only a selection of eigenstates is shown here. The full figure up to $J\leq 2$ can be found in [35], but it provides the same qualitative picture.
Figure 1: Allowed dipole (red squares) and quadrupole transitions (blue squares) between eigenstates of the asymmetric top Hamiltonian. Panel columns represent different circular polarizations of the beam, and panel rows – the amount of the OAM that can be transferred from the beam to the internal rotation of the molecule. Purple circles show an example of spin-orbit coupling creating an $E_{1}$-$E_{2}$ term with the transfer of OAM. Orange squares highlight the transitions that would be superimposed in the experiment, as shown in Fig. 2.
The three columns of Fig. 1 correspond to different beam polarizations, and the three rows to different OAM transfer quantum numbers $m$. As shown in Ref. [23, 24], electric dipole transitions do not transfer OAM, so they appear only in the $m=0$ row (though we display them in the $m=\pm 1$ rows for comparison). In contrast, electric quadrupole transitions can transfer up to one OAM quantum ($m=0,\pm 1$). Any non-planar beam profile generally contains contributions in all three rows, but their relative weights depend on the specific beam structure.
@@
However, the clear separation of the OAM in terms of $m$ and the spin in terms of $\sigma$ is only valid for paraxial beams. In fact, any kind of SOC of light would result in a coupling of different $\sigma$ components [25, 37]. For our purposes, this means that the terms from different panels in Fig. 1 are enabled to drive the same transition. This would, for example, give rise to the $E_{1}$-$E_{2}$ term corresponding to the purple square with a transition that is simultaneously dipole (at $m=0$, $\sigma=-1$) and quadrupole (at both $m=0$, $\sigma=-1$ and $m=1$, $\sigma=0$): $|1_{0,1}(1)\rangle\to|1_{1,1}(0)\rangle$. The coupling of the dipole moment term to $m=-1$ quadrupole component enables the molecule to sense the impact of OAM.
The spin-orbit coupling of light, required to realize such transitions, can be achieved by tightly–focused Laguerre-Gaussian beams [17, 38, 39]. Here, we describe the beam in the first–order non-paraxial approximation [40] (for details see [35]), and use the interaction Hamiltonian of Ref. [23, 24] to calculate the spatially-resolved matrix elements between the states denoted by a dashed square in Fig. 1. The chosen parameters are transition quadrupole moment $Q_{xy}=0.5d_{z}\lambda$, where $d_{z}$ is the transition dipole moment and $\lambda$ the wavelength, focal beam-waist $\omega_{0}=0.2\lambda$, distance from focus $z=\lambda$, with $z$ being the propagation direction, azimuthal charge of the Laguerre-Gausian $L_{\rm beam}=1$ and polarization $\sigma=-1$. The opposite enantiomers are described by reflection on the molecule fixed (inertial) $xy$-plane. It is well-enstablished [41, 42] that an anisotropy is needed to observe the $E_{1}$-$E_{2}$  interaction, which is why $M$ state resolution is required in any proposed HD experiment. The obtained signal profiles for states with resolved $M$ are shown in Fig. 2.
Figure 2: Signal profiles of the $E_{1}$-$E_{2}$ transitions from $\bra{1_{1,1}(M)}$ | to |$\ket{1_{0,1}(M^{\prime})}$ for the opposite enantiomers coupled by a Laguerre Gaussian beam without (top) and with (middle row) the spin-orbit coupling and by a Gaussian beam with spin-orbit coupling (bottom).
We first note that the integral of all shown profiles is equal, unless $d_{z}Q_{xy}^{*}$ has an imaginary component ${\rm Im}[d_{z}Q_{xy}^{*}]$, which leads to dichroism on the level of total absorption. This is a known source of vibrational CD, and is independent of the structure of the electric field and only depends on the internal parameters of the molecule, see [41]. From this result we conclude that HD is impossible to detect by just the magnitude of the absorbed intensity of light, but requires the spatial resolution of the absorption profile [17].
@@
The focal field introduces SOC of light, which gives rise to new distinguishing effects between the enantiomers in the absorption profiles, seen in the middle row of Fig. 2. First, there is a large difference in the absorption of $M=-1\to M^{\prime}=0$ and $M=0\to M^{\prime}=1$ transitions. The signal comes from coupling the dipole and quadrupole transitions in the dashed squares in Fig. 1. This difference stems from the relative handedness of the vortex beam and the chiral molecule and involves transfer of OAM. Therefore, this can be considered genuine HD.
@@
Next, there is also a difference in the $M=1\to M^{\prime}=1$ and $M=-1\to M^{\prime}=-1$ profiles. This arises from the coupling of dipole transitions in the dashed squares with $\sigma=0$ (Fig. 1) with the quadrupole transitions in the $m=-1$, $\sigma=-1$ square. Therefore, this is also HD. Overall, the dichroism signal in the case of SOC light is an order of magnitude stronger than without the coupling.
@@
The bottom row in Fig. 2 shows that in the case of SOC, it is not necessary to use a beam with OAM in the asymptotic limit. Indeed, HD is also observed in the case of a tight focused Gaussian beam. This is because the focal field takes the form of LG mode of azimuthal charge $\sigma$. In fact, a similar dichroism is also seen for a tight focused linearly polarized beam, because the two components of circular polarization within it couple through the SOC. This shows that neither OAM nor spin of the incoming beam are an essential component to observe HD, but rather the SOC.
@@
Note, that due to the SOC of light the type of angular momentum transferred to the molecule cannot be easily distinguished which hinders the characterization of the resulting dichroism as either CD or HD. Our work pinpoints the OAM and spin absorption channels in terms of the quantum numbers $m$ and $\sigma$ (see Fig. 2) providing a way to resolve this issue.
@@
In summary, we have shown that paraxial OAM beams cannot produce HD, as any observed dichroism under these conditions reduces to standard CD. Genuine HD – dichroism directly tied to the transfer of OAM – requires spin–orbit coupling of light, as in tightly focused beams. Strikingly, a purely Gaussian mode can exhibit genuine HD once sufficiently focused, even though it carries no OAM in the far-field. These findings clarify the origin of HD in structured-light experiments and suggest that prior measurements of HD may warrant a careful re-examination to confirm the presence of OAM transfer. In particular, the studies that did not observe dichroism were either probing primarily dipole-order interactions [16] or lost the dichroism signal due to averaging [17]. The studies that successfully detected dichroism, on the other hand, employed both quadrupole field interactions and tightly focused beams [18, 20, 21, 22].
The framework developed here offers a roadmap for designing chiral-sensitive spectroscopy and enantiomer separation methods leveraging the spatial structure of light [43].

### v2

@@ Abstract
We present a general theoretical framework for helical dichroism (HD), establishing an explicit link between chiral resolution and orbital angular momentum (OAM) exchange in light–matter interaction. Tracing microscopic mechanisms of the OAM transfer, we derive rotational selection rules, which establish that HD emerges only from the spin–orbit coupling of light, even for beams without the far-field OAM. Our findings refine the conditions for observing HD, provide a tool to re-examine the outcome of prior experiments, and guide future designs for chiral sensing with structured light.
@@
For decades, the difference in absorption of left- and right-circularly polarised light, known as the circular dichroism (CD) [6, 7], has been used in spectroscopy of chiral media [8]. However, circular dichroism relies on a limited resource, namely, the spin angular momentum of light. Meanwhile, the pioneering study by Allen et al. [9] has demonstrated that, in addition to spin, light can also carry orbital angular momentum (OAM). In contrast to spin, the OAM of a beam is, in principle, unbounded. The helical wavefronts of twisted light, also known as optical vortex beams, fascinated physicists, who envisaged that their chiral structure could match the chirality of matter, offering unprecedented possibilities for enantioselective applications [10].
@@
This fascination culminated in various concrete OAM-induced processes being proposed as probes of enantiomeric purity, particularly, in the form of helical dichroism (HD) [11, 12, 13], which is argued to represent a significant improvement over CD [12, 13]. Theoretical works of Forbes and Andrews [14, 15, 16] laid the groundwork for HD, by carefully analyzing the symmetry properties of the terms in the multipolar expansion of light-matter interaction, and pinpointing the electric-dipole–magnetic-dipole ($E_{1}$-$M_{1}$), electric-dipole–electric-quadrupole ($E_{1}$-$E_{2}$) and electric-quadrupole–electric-quadrupole ($E_{2}$-$E_{2}$) terms as chiral discriminators. These authors derived the expression for the $E_{1}$-$E_{2}$ transition amplitude in terms of molecular (transition) multipole moments; however, they made no attempt to evaluate the latter in terms of molecular rotational states. On the experimental side, initial attempts showed no signs of HD [17, 18]; however, recent experiments have claimed to observe it [19, 20, 21, 22, 23]. Nonetheless, the experimental apparatus of these experiments is very complex, and, although it was argued that tight focusing of a beam is a requirement [15], so far there have been no theoretical attempts to pinpoint the origin of the observed dichroism.
Clearly, a microscopic theory is needed that (i) links the observed dichroic signals to contributing ro-vibrational transitions and the associated transfer of different types of angular momenta, and (ii) identifies the required molecular symmetries or focusing conditions that enable HD. This work fills this gap in the literature by establishing an explicit connection between chiral resolution and orbital angular momentum exchange in light-molecule interaction, and identifying the conditions that give rise to HD in ro-vibrational spectroscopy, paving the way for further advancements of this field.
@@
We first provide an overview of chirality detection methods based on molecular multipole moments, and derive symmetry requirements for observing $E_{1}$-$E_{2}$-based dichroism. Using our recently introduced molecule-light interaction Hamiltonian [24, 25], we derive the selection rules for ro-vibrational transitions induced by the $E_{1}$-$E_{2}$ term. By tracing the angular momentum transfer involved in these rules, we demonstrate that paraxial Laguerre-Gaussian (LG) beams cannot produce the genuine HD, i.e., HD involving the transfer of OAM from light to the molecule. Instead, we show that tight-focusing associated with the spin-orbit coupling (SOC) of light [26] enables OAM transfer to chiral molecules by introducing HD contributions that are absent within the paraxial approximation. Strikingly, even a Gaussian beam, which carries no OAM in the far-field, can produce a HD signal in the tightly-focused regime. Collectively, our findings refine the theoretical foundations of HD in ro-vibrational spectroscopy, and guide the designs of future experiments and enantiomer-sensitive techniques using structured light.
@@
Existing enantiomer-sensitive spectroscopy techniques rely on the analysis of the multipole moments of the molecule. The (vibrational) circular dichroism—the standard method to detect chirality—is based on the $E_{1}$-$M_{1}$ coupling and has already been exhaustively investigated [27, 28, 29, 30]. It draws on the fact that the electric dipole moment of a chiral molecule transforms as a vector, and thus picks up a minus sign, under mirror reflection, while the magnetic dipole, which is a pseudovector, is invariant to it. A coexistence of both these multipoles implies that a molecule lacks a mirror symmetry plane (strictly speaking, an improper rotation axis) and is, by definition, chiral. This type of CD can be exhibited by chiral molecules, for which the symmetry point group permits at least one nonzero dipole moment. However, the dichroism stemming from the $E_{1}$-$M_{1}$ term only depends on the intensity of the field, which makes other sources, where the dichroism can be enhanced by the structure of the field, attractive alternatives.
@@
One example—the phase sensitive microwave three-wave mixing technique [31, 32]—exploits the fact that a non-zero product of all dipole moments $d_{x}d_{y}d_{z}\neq 0$ implies that the molecule is chiral [32, 33, 34]. Symmetry analysis reveals that only completely asymmetric ($C_{1}$) chiral molecules can be characterized by this condition.
@@
Our work explores another direction – the conditions for chirality stemming from the electric dipole and quadrupole moments of the molecule. The dichroism induced by the $E_{1}$-$E_{2}$ term can be enhanced with appropriately structured fields, which makes it a promising prospect for future enantiomer-sensitive techniques. Recently, the magnetic-dipole-magnetic quadrupole ($M_{1}$-$M_{2}$) coupling has also been suggested as a source of HD, which is equivalent to the $E_{1}$-$E_{2}$ term from a symmetry standpoint [35]. However, this and other higher-order terms, such as $E_{2}$-$E_{2}$, are difficult to detect experimentally due to low transition probabilities in the multipole expansion.
@@
The remainder of this paper focuses on characterizing the dichroism originating from the $E_{1}$-$E_{2}$ term. Using a simple point-charge model, we derive the following condition (proof in the Supplementary Material  [^1] See Supplementary Material, containing references [24, 25, 41, 26, 38]):
@@
If the product of dipole and quadrupole moments $d_{\mu}Q_{\nu\rho}\neq 0$ for distinct coordinates $\mu,\nu,\rho\in\{x,y,z\}$ of the principal axis system, then the molecule is chiral.
@@
To date, the above condition has always been associated with HD [11, 14]. However, as we show in this work, it can also lead to discriminatory effects without an OAM transfer from light to the internal molecular degrees of freedom. To resolve this ambiguity, this work uses “helical dichroism” to denote dichroism effects where OAM transfer occurs, while dichroism without such transfer, even for an OAM-carrying beam, is referred to as CD.
@@
Table 1 summarizes which chiral point groups can exhibit $E_{1}$-$E_{2}$-induced dichroism. In the case of pure-rotational spectroscopy, only the permanent multipole moments of the molecule are relevant. In ro-vibrational spectroscopy vibrational transitions can alter the symmetry of a molecular state. Therefore, transformation properties of vibrational states in terms of their associated irreducible representation $V$ need to be considered, which allows for the detection of more chiral groups. Table 1 provides the list of $V$’s exhibiting dichroism for each chiral point group.
@@
Table 1: The $E_{1}$-$E_{2}$ chiral signatures in rotational and ro-vibrational spectroscopy of all possible chiral point groups. In the case of ro-vibrational spectroscopy, we provide the irreducible-representations of the vibrational modes associated with non-zero chiral signature.
Point group $C_{1}$ $C_{2}$ $C_{n\geq 3}$ $D_{2}$ $D_{n\geq 3}$ $T$ $O$ II
@@
Ro-vibrational AA AA, $B$ $E_{1}$ $B_{1,2,3}$ $E_{1}$ $T$
As displayed in Table 1 the $E_{1}$-$E_{2}$ term can only identify chiral molecules with $C_{1}$ and $C_{2}$ symmetries via transitions that preserve the molecular symmetry (pure rotational and ro-vibrational of the trivial representation $A$). Since most known chiral molecules are $C_{1}$-symmetric, i.e., they are asymmetric tops [37], we restrict our scope to these molecules.
[…]
Here, $J$ is total angular momentum quantum number, $M$ and $K$ are the projections of $\bm{J}$} on the lab quantization axis $z$, herewith fixed to the propagation axis of light, and the body-fixed principal axis respectively. The asymetric top rotor indices $N_{p}$ and $N_{o}$ are the limits of $K$ in the pure prolate/oblate solutions and are not good quantum numbers. Since quantum numbers $J$ and $M$ are not affected by the change in basis (1), we proceed to derive the selection rules for the electric dipole ($E_{1}$) and quadrupole ($E_{2}$) transitions within the rotational basis, $|J,M,K\rangle$.
[…]
@@
where $l=0$ and $l=1$ for dipole and quadrupole transitions respectively, $\mu=0,\pm 1,\dots,\pm(l+1)$ runs over all components of a given multipole moment in the molecule-fixed frame, while both $m=0,\pm 1,\dots,\pm l$, the OAM transferred to the molecule, and $\sigma=0,\pm 1$, the polarization component of light, correspond to lab-frame angular momenta. The vibrational, $\mathcal{I}^{\text{vib};l,\mu}$, and center-of-mass, $\mathcal{I}^{\text{CM};l,m}$, terms do not depend on molecular rotational states. Hence, in this work, we focus on the rotational transition amplitude $\mathcal{I}^{\text{rot};l,m,\mu,\sigma}_{J,M,K,J,M^{\prime},K^{\prime}}$, which is proportional to the Clebsh-Gordan coefficient $C^{J,M}_{j_{1},m_{1},j_{2},m_{2}}$ [24, 25]. From this coefficient, we derive the angular momentum selection rules, summarized in Table 2 [36].
[…]
@@
$|J,0,K\ket$$\not\to\ket{J,0,K^{\prime}}$ for details)
To realize an $E_{1}$-$E_{2}$ transition, both $E_{1}$ ($l=m=0$) and $E_{2}$ ($l=1$, $m=0,\pm 1$) transitions should be allowed between the chosen two states simultaneously. According to the $\Delta M$ selection rules (see Table 2), for light with a well-defined polarization, $\sigma$, this is only possible for the $m=0$ component of the beam, which carries no OAM. This brings us to the first key finding of our work: The dichroism stemming from the $E_{1}$-$E_{2}$ term of a paraxial beam should not be associated with HD, as no OAM transfer to the molecule occurs. Moreover, this dichroism is present even in the case of a plane-wave and thus is not associated at all to the structure of light or its helicity.
This raises the question whether the $E_{1}$-$E_{2}$ term can produce HD, where OAM transfer occurs. In what follows, we show that we can create such a scenario in the case of non-paraxial beams.
In Fig. 1, we demonstrate the allowed dipole (red squares) and quadrupole (blue squares) transitions between the rotational eigenstates of asymmetric top molecules. For clarity, only a selection of eigenstates is shown here. The full figure up to $J\leq 2$ can be found in Ref. [36], but it provides the same qualitative picture.
Figure 1: Allowed dipole (red squares) and quadrupole transitions (blue squares) between eigenstates of the asymmetric top Hamiltonian. Panel columns represent different circular polarizations of the beam, and panel rows – the amount of the OAM that can be transferred from the beam to the internal rotation of the molecule. Purple circles show an example of SOC creating an $E_{1}$-$E_{2}$ term with the transfer of OAM. Orange squares highlight the transitions that would be superimposed in the experiment, as shown in Fig. 2.
The three columns of Fig. 1 correspond to different beam polarizations $\sigma$, and the three rows to different OAM transfer quantum numbers $m$. As shown in Ref. [24, 25], electric dipole transitions do not involve the OAM transfer, so they appear only in the $m=0$ row (though we also display them in the $m=\pm 1$ rows for comparison). In contrast, electric quadrupole transitions can involve the transfer of up to one OAM quantum ($m=0,\pm 1$). Any non-planar beam profile generally contains contributions in all three rows, but their relative weights depend on the specific beam structure.
@@
However, the clear separation of the OAM in terms of $m$, and spin in terms of $\sigma$, is only valid for paraxial beams. In fact, any kind of SOC of light would result in a coupling of different $\sigma$ components [26, 38]. For our purposes, this means that the terms from different columns in Fig. 1 are enabled to drive the same transition. This would, for example, give rise to the $E_{1}$-$E_{2}$ term corresponding to the purple square with a transition that is simultaneously dipole (at $m=0$, $\sigma=-1$) and quadrupole (at both $m=0$, $\sigma=-1$ and $m=1$, $\sigma=0$): $|1_{0,1}(1)\rangle\to|1_{1,1}(0)\rangle$. The coupling of the dipole moment term to the quadrupole component with m=1m=-1 enables the molecule to absorb the OAM of light since $m\neq 0$.
The SOC of light, required to realize such transitions, can be achieved, e.g., by the tight focusing of Laguerre-Gaussian beams [18, 39, 40]. The electric field at the focus of the beam (focal field) induces the interconversion between the spin and OAM, manifesting as the SOC of light [36, 26, 38]. Here, we describe the beam in the first-order non-paraxial approximation [36, 41], and use the interaction Hamiltonian of Ref. [24, 25] to calculate the spatially-resolved matrix elements between the states denoted by a dashed square in Fig. 1. The chosen parameters are: transition quadrupole moment $Q_{xy}=0.5\,d_{z}\lambda$, where $d_{z}$ is the transition dipole moment and $\lambda$ the wavelength, focal beam-waist $\omega_{0}=0.2\,\lambda$, distance from focus $z=\lambda$, with $z$ being the propagation direction, azimuthal charge of the Laguerre-Gausian $L_{\rm beam}=1$ and polarization $\sigma=-1$. Note that changing $L_{\rm beam}\to-L_{\rm beam}$ would only affect the relative amplitudes of the $M^{\prime}=-1\to M=1$ and $M^{\prime}=M=0,\pm 1$ transitions, and not the spatial distributions. The opposite enantiomers are described by reflection on the molecule fixed (inertial) $xy$-plane. It is well-enstablished [42, 43] that an anisotropy is needed to observe the $E_{1}$-$E_{2}$  interaction, which is why $M$ state resolution is required in any proposed HD experiment. The obtained signal profiles for states with resolved $M$ are shown in Fig. 2.
[…]
We first note that the integral of all shown profiles is equal, unless $d_{z}Q_{xy}^{*}$ has an imaginary component ${\rm Im}[d_{z}Q_{xy}^{*}]$, which leads to dichroism at the level of total absorption. This is a known source of vibrational CD, and is independent of the structure of the electric field and only depends on the internal parameters of the molecule, see [42]. From this result we conclude that HD is impossible to detect by just the magnitude of the absorbed intensity of light, but requires the spatial resolution of the absorption profile [18].
@@
Tight focusing introduces SOC of light, which gives rise to new distinguishing effects between the enantiomers in the absorption profiles, seen in the middle row of Fig. 2. First, there is a large difference in the spatially-resolved absorption of $M=-1\to M^{\prime}=0$ and $M=0\to M^{\prime}=1$ transitions. The signal comes from the coupling of dipole and quadrupole transitions in the dashed squares in Fig. 1. This difference stems from the relative handedness of the vortex beam and the chiral molecule and involves transfer of OAM. Therefore, this can be considered genuine HD.
@@
Next, there is also a difference in the $M=1\to M^{\prime}=1$ and $M=-1\to M^{\prime}=-1$ profiles. It arises from the coupling of dipole transitions in the dashed squares with $\sigma=0$ (Fig. 1) to the quadrupole transitions in the m=1m=-1, $\sigma=-1$ square. Therefore, this is also HD. Overall, the dichroism signal in the case of SOC light is an order of magnitude stronger than without the coupling.
@@
The bottom row in Fig. 2 shows that in the case of SOC, it is not necessary to use a beam with OAM in the asymptotic limit. Indeed, HD is also observed in the case of a tightly-focused Gaussian beam. This is because the focal field, through the conversion of spin into OAM, takes the form of an LG mode with the azimuthal charge $\sigma$. In fact, a similar dichroism is also seen for a tightly-focused linearly polarized beam, because the two components of circular polarization within it couple through the SOC. This shows that neither OAM nor spin of the incoming beam are an essential component to observe HD, but rather the SOC.
@@
Note, that due to the SOC of light, the type of angular momentum transferred to the molecule cannot be easily distinguished, which hinders the characterization of the resulting dichroism as either CD or HD. Our work pinpoints the OAM and spin absorption channels in terms of the quantum numbers $m$ and $\sigma$ (see Fig. 2) providing a way to resolve this issue.
@@
In summary, we have shown that paraxial vortex beams cannot produce HD, as any observed dichroism under these conditions reduces to the standard CD. Genuine HD—dichroism directly tied to the transfer of OAM—requires SOC of light, as in the case of tightly-focused beams. Strikingly, a purely Gaussian mode can exhibit genuine HD once sufficiently focused, even though it carries no OAM in the far-field. These findings clarify the origin of HD in structured-light experiments and suggest that prior measurements of HD may warrant a careful re-examination to confirm the presence of OAM transfer. In particular, the studies that did not observe dichroism were either probing primarily dipole-order interactions [17] or suppressed the dichroism signal due to averaging [18]. The studies that successfully detected dichroism, on the other hand, employed both quadrupole field interactions and tightly-focused beams [19, 21, 22, 23]. To properly model the full complexity of these experiments, the here-presented theory should in the future be expanded to include the electronic degrees of freedom. This will allow us to address processes that involve electronic transitions, including complex multi-photon processes, and explain new emerging experiments in both atoms and molecules [44].
In light of our results, the experimental setup for ro-vibrational chiral-sensitive spectroscopy and enantiomer separation that leverages the spatial structure of light is expected to be complex and demanding [45]. Nevertheless, the framework developed here offers a roadmap for future experimental designs and a versatile tool to quantify both HD and the newly identified CD signatures.